# Robust Neural Information Retrieval: An Adversarial and Out-of-distribution Perspective

### v1

@@ 5.1.2. Challenges on OOD robustness in IR
Many key issues in OOD robustness in IR have not received much attention.
@@
The reliance on large-scale data sets for training and evaluating IR systems poses significant challenges in ensuring OOD robustness. Large datasets often contain biases and do not adequately represent the diversity of real-world scenarios (Kaplan et al., 2014), leading to models that perform well on seen data but poorly on unseen, OOD examples (Thakur et al., 2021; Yu et al., 2022). Addressing this challenge requires innovative approaches to data collection and model training that prioritize diversity and real-world applicability, ensuring that IR systems remain reliable and effective across a broad range of OOD scenarios.
@@
Existing work mainly considers the problem of one-shot adaptation of DRMs to new corpora (Thakur et al., 2021; Cohen et al., 2018). However, in real search engines, new documents are constantly added to the search engine and bring in a variety of new domains, which poses a challenge to the continuous adaptation ability of DRMs. To address this, it is essential to develop DRMs that not only quickly adapt to new corpora in a one-shot learning scenario but also continuously learn and adjust as new data is introduced. This requires innovative approaches that can dynamically update the models in an incremental fashion without the need for frequent retraining from scratch. Techniques such as online learning, transfer learning, and meta-learning can play pivotal roles in enhancing the continuous corpus adaptation of DRMs, ensuring they remain effective and relevant in the ever-changing search landscape.
@@
Existing approaches to improve OOD generalization ability mainly target DRMs (Thakur et al., 2021; Yu et al., 2022; Zhan et al., 2022). As an important part of the search engine, the OOD robustness of the NRMs should be optimized along with the DRMs. Enhancing the OOD generalization of NRMs involves developing models that can effectively handle queries that deviate significantly from the training distribution, thereby ensuring the retrieval of relevant and accurate results under a wide range of search scenarios. This challenge requires a multifaceted approach, incorporating advanced machine learning techniques such as robust representation learning, anomaly detection, and domain adaptation strategies. By prioritizing the OOD robustness of NRMs alongside DRMs, search engines can significantly improve their ability to serve high-quality, relevant content to users, even when faced with novel or unexpected queries.
Improving OOD robustness in practical search engines
Enhancing OOD robustness in practical search engines is essential for maintaining their effectiveness, efficiency, and cost-effectiveness. To achieve this, search engines must be able to accurately identify and process queries that fall outside the typical distribution of observed data, ensuring that even uncommon or unseen queries return relevant and useful results. Implementing robust OOD handling mechanisms can significantly improve the quality of search results (Thakur et al., 2021; Liu et al., 2023b), but this often requires sophisticated algorithms that can detect and adapt to OOD queries in real-time (Zhuang and Zuccon, 2022; Penha et al., 2022). While these algorithms can be computationally intensive, leading to higher operational costs, the investment in OOD robustness can ultimately enhance user satisfaction and trust in the search engine (Liu et al., 2023b; Zhuang et al., 2023). Moreover, optimizing these algorithms for efficiency can help mitigate additional costs, ensuring that improvements in OOD robustness also align with the search engine’s operational goals.
@@ 5.2. Challenges and Opportunities Posed by Large Language Models
Recently, LLMs have gained attention for their excellent generative abilities. There have been several works applying them to IR with good results. We believe that the introduction of LLMs may bring new robustness problems and also provide new solutions to existing robustness problems.
@@ 5.2.1. New challenges to IR robustness from LLMs
Recently, there has been a lot of beginning work on using LLMs for IR tasks. However, these attempts may pose new robustness challenges to LLMs-based IR methods due to the robustness issues of LLMs themselves.
@@
When applied to IR systems, the adversarial vulnerability of the LLMs themselves is introduced at the same time, as demonstrated by:
  • The vulnerability caused by hallucinations of LLMs. There has been some work using the LLM as a ranking model (Sun et al., 2023). However, LLMs can generate plausible yet factually incorrect or irrelevant information—a phenomenon known as “hallucination”. Existing work has developed attacks against LLM-based ranking models, and vulnerability issues have been found with such models (Liu et al., 2024a). In IR systems, such reliance can lead to the retrieval of misleading or inaccurate data, undermining the trustworthiness and reliability of the system. Addressing these hallucinations is crucial for maintaining the integrity of search results and the credibility of IR systems using LLMs (Shuster et al., 2021; Wu et al., 2023b).
  • Defense costs associated with the scale and opacity of LLMs. LLMs are not only large and complex but also often operate as black boxes with limited transparency into how decisions are made. This opacity complicates efforts to diagnose and mitigate vulnerabilities, making defensive measures both technically challenging and resource-intensive. As LLMs scale, the cost and complexity of implementing robust defensive strategies increase exponentially.
@@
LLMs have shown biases and input sensitivities in existing work (Ji et al., 2023; Ferrara, 2023), and these will affect the OOD robustness of IR systems:
@@
  • Sensitivity of LLMs to query inputs. LLMs can exhibit high sensitivity to slight variations in input (Ji et al., 2023; Ferrara, 2023), potentially leading to significantly different IR outcomes. This sensitivity can be particularly problematic in IR contexts where consistent and predictable retrieval performance is necessary. Enhancing the robustness of LLMs against input variation is a critical area for research and development.
@@ 5.2.2. New opportunities for IR robustness via LLMs
While the use of LLMs may introduce new robustness risks, the power of LLMs also provides new ideas for improving robustness. In the field of NLP, many works are used to enhance the robustness of NLP models using LLMs (Moraffah et al., 2024; Helbling et al., 2023; Zeng et al., 2024), but not yet in IR.
@@
  • Synthesizing OOD training data with LLMs. LLMs can generate diverse and complex datasets that mirror OOD scenarios (Bonifacio et al., 2022), providing IR systems with the training needed to better handle unfamiliar or novel situations. Such synthetic data can help improve the generalizability and robustness of IR models against OOD inputs.
@@ 6. Conclusion
In conclusion, robust neural IR represents a complex and multifaceted problem, but also an opportunity for innovative research and development. As we look forward, the goal of developing IR systems that are not only robust but also adaptable, ethical, and user-centric is essential, promising to redefine the boundaries of what is possible in information retrieval.
@@ Acknowledgements.
This work was funded by the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, the National Key Research and Development Program of China under Grants No. 2023YFA1011602 and 2021QY1701, the National Natural Science Foundation of China (NSFC) under Grants No. 62372431, the Youth Innovation Promotion Association CAS under Grants No. 2021100, the Lenovo-CAS Joint Lab Youth Scientist Project, and the project under Grants No. JCKY2022130C039. This work was also (partially) funded by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), project ROBUST with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO), DPG Media, RTL, and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023, and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.

### v2

@@ 5.1.2. Challenges on OOD robustness in IR
Several key issues in OOD robustness in IR have not received much attention.
@@
The reliance on large-scale data sets for training and evaluating IR systems poses significant challenges in ensuring OOD robustness. Large datasets often contain biases and do not necessarily represent the diversity of real-world scenarios (Kaplan et al., 2014), leading to models that perform well on seen data but poorly on unseen, OOD examples (Thakur et al., 2021; Yu et al., 2022). Addressing this challenge requires innovative approaches to data collection and model training that prioritize diversity and real-world applicability, ensuring that IR systems remain reliable and effective across a broad range of OOD scenarios.
@@
As we have seen in Section 3 and 4, prior work mainly considers the problem of one-shot adaptation of DRMs to new corpora (Thakur et al., 2021; Cohen et al., 2018). However, in real search engines, new documents are constantly added to the search engine and bring in a variety of new domains, which poses a challenge to the continuous adaptation ability of DRMs. To address this, it is essential to develop DRMs that not only quickly adapt to new corpora in a one-shot learning scenario but also continuously learn and adjust as new data is introduced. This requires innovative approaches that can dynamically update the models in an incremental fashion without the need for frequent retraining from scratch. Techniques such as online learning, transfer learning, and meta-learning can play pivotal roles in enhancing the continuous corpus adaptation of DRMs, ensuring they remain effective and relevant in the ever-changing search landscape.
@@
Approaches to improve OOD generalization ability mainly target neural retrieval models (Thakur et al., 2021; Yu et al., 2022; Zhan et al., 2022). More generally, the OOD robustness of neural ranking models should be optimized along with retrieval models. Enhancing the OOD generalization of NRMs involves developing models that can effectively handle queries that deviate significantly from the training distribution, thereby ensuring the retrieval of relevant and accurate results under a wide range of search scenarios. This challenge requires a multifaceted approach, incorporating advanced machine learning techniques such as robust representation learning, anomaly detection, and domain adaptation strategies. By prioritizing the OOD robustness of NRMs alongside DRMs, search engines can significantly improve their ability to serve high-quality, relevant content to users, even when faced with novel or unexpected queries.
Improving OOD robustness in practice
Enhancing OOD robustness in practice is essential for maintaining the effectiveness, efficiency, and cost-effectiveness of search engines. To achieve this, search engines must be able to accurately identify and process queries that fall outside the typical distribution of observed data, ensuring that even uncommon or unseen queries return relevant and useful results. Implementing robust OOD handling mechanisms can significantly improve the quality of search results (Thakur et al., 2021; Liu et al., 2023b), but this often requires sophisticated algorithms that can detect and adapt to OOD queries in real-time (Zhuang and Zuccon, 2022; Penha et al., 2022). While these algorithms can be computationally intensive, leading to higher operational costs, the investment in OOD robustness can ultimately enhance user satisfaction and trust in the search engine (Liu et al., 2023b; Zhuang et al., 2023). Moreover, optimizing these algorithms for efficiency can help mitigate additional costs, ensuring that improvements in OOD robustness also align with the search engine’s operational goals.
@@ 5.2. Challenges and Opportunities Posed by Large Language Models
LLMs have gained attention for their generative abilities. There have been several works applying them to IR with good results. The introduction of LLMs may bring new robustness problems and also provide new solutions to known robustness problems.
@@ 5.2.1. New challenges to IR robustness from LLMs
Recently, there has been a lot of exploratory work on using LLMs for IR tasks. However, these attempts may pose new robustness challenges to LLMs-based IR methods due to the robustness issues of LLMs themselves.
@@
When applied to IR systems, the adversarial vulnerability of the LLMs themselves is imported at the same time, as demonstrated by:
  • The vulnerability caused by hallucinations of LLMs. There has been some work using the LLM as a ranking model (Sun et al., 2023). However, LLMs can generate plausible yet factually incorrect or irrelevant information, a phenomenon known as “hallucination.” Prior work has developed attacks against LLM-based ranking models, and vulnerability issues have been found with such models (Liu et al., 2024a). In IR systems, such reliance can lead to the retrieval of misleading or inaccurate data, undermining the trustworthiness and reliability of the system. Addressing these hallucinations is crucial for maintaining the integrity of search results and the credibility of IR systems using LLMs (Shuster et al., 2021; Wu et al., 2023b).
  • Defense costs associated with the scale and opacity of LLMs. LLMs are not only large and complex but also often operate as black boxes with limited transparency into how decisions are made. This opacity complicates efforts to diagnose and mitigate vulnerabilities, making defensive measures both technically challenging and resource-intensive. As LLMs scale, the costs and complexity of implementing robust defensive strategies are likely to increase substantially.
@@
LLMs have shown biases and input sensitivities (Ji et al., 2023; Ferrara, 2023), and these will affect the OOD robustness of IR systems:
@@
  • Sensitivity of LLMs to query inputs. LLMs can exhibit high sensitivity to slight variations in input (Chen et al., 2024; Ji et al., 2023; Ferrara, 2023), potentially leading to significantly different IR outcomes. This sensitivity can be particularly problematic in IR contexts where consistent and predictable retrieval performance is necessary. Enhancing the robustness of LLMs against input variation is a critical area for research and development.
@@ 5.2.2. New opportunities for IR robustness via LLMs
While the use of LLMs may introduce new robustness risks, the power of LLMs also provides new ideas for improving robustness. In the field of NLP, many works are used to enhance the robustness of NLP models using LLMs (Moraffah et al., 2024; Helbling et al., 2023; Zeng et al., 2024), but not so much yet in IR.
@@
  • Synthesizing OOD training data with LLMs. LLMs can generate diverse and complex datasets that mirror OOD scenarios (Bonifacio et al., 2022; Rajapakse and de Rijke, 2023), providing IR systems with the training needed to better handle unfamiliar or novel situations. Such synthetic data can help improve the generalizability and robustness of IR models against OOD inputs.
@@ 6. Conclusion
In conclusion, robust neural IR represents a complex and multifaceted problem space, but also an opportunity for innovative research and development. As we look forward, the goal of developing IR systems that are not only robust but also adaptable, trustworthy, and user-centric is essential, promising to redefine the boundaries of what is possible in information retrieval.
@@ Acknowledgements.
This work was funded by the National Key Research and Development Program of China under Grants No. 2023YFA1011602 and 2021QY1701, the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, the National Natural Science Foundation of China (NSFC) under Grants No. 62372431, the Youth Innovation Promotion Association CAS under Grants No. 2021100, the Lenovo-CAS Joint Lab Youth Scientist Project, and the project under Grants No. JCKY2022130C039. This work was also (partially) funded by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), project ROBUST with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO), DPG Media, RTL, and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023, and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.
Appendix A Source Selection
We loosely followed the guidelines by Kitchenham and Charters (2007) for the selection of publications used in our survey.
A.1. Sources
We used the sources listed in Table 4 for our survey.
Table 4. Venues, journals, and repositories used for the survey.
Source Acronym or URL
AAAI Conference on Artificial Intelligence AAAI
Annual Meeting of the Association for Computational Linguistics ACL
arXiv https://arxiv.org
ACM Conference on Computer and Communications Security CCS
ACM International Conference on Information and Knowledge Management CIKM
European Conference on Information Retrieval ECIR
Conference on Empirical Methods in Natural Language Processing EMNLP
Foundations and Trends in Information Retrieval FnTIR
International Conference on Learning Representations ICLR
International Conference on Data Mining ICDM
International Conference on Machine Learning ICML
International Conference on the Theory of Information Retrieval ICTIR
International Joint Conference on Artificial Intelligence IJCAI
Information Processing & Management IPM
ACM SIGKDD Conference on Knowledge Discovery and Data Mining KDD
Annual Conference of the North American Chapter of the Association for Computational Linguistics NAACL
Conference on Neural Information Processing Systems NeurIPS
SIGIR Conference on Research and Development in Information Retrieval SIGIR
Transactions of the Association for Computational Linguistics TACL
ACM Transactions on Information Retrieval TOIS
Text Retrieval Conference TREC
International World Wide Web Conference WebConf
A.2. Inclusion criteria
For papers in the sources listed in Table 4 we used the following criteria to include them in our survey:
  1. (IC1)
    The paper proposes a definition of one or several robustness notions in the context of IR.
  2. (IC2)
    The paper proposes an approach to improving the robustness of an IR model.
  3. (IC3)
    The paper proposes a method to evaluate one or several robustness notions in the context of IR.
  4. (IC4)
    The paper presents one or several benchmarks for assessing the robustness of IR models.
  5. (IC5)
    The paper presents a study that investigates the foundations of robustness of IR models.
A.3. Exclusion criteria
For papers in the sources listed in Table 4 we used the following criteria to exclude them in our survey:
  1. (EC1)
    The paper is not written in English.
  2. (EC2)
    The paper is not in the date range of January 2012 to July 2024.
  3. (EC3)
    An extended version of the paper has been published, which subsumes its contents.
Table 5. Statistics of datasets in BestIR benchmark. #Doc denotes the number of documents in corpus; #Qtrain denotes the number of queries available for training; #Qdev denotes the number of queries available for development; #Qeval denotes the number of queries available for evaluation.
Robustness Type Dataset #Doc #Qtrain #Qdev #Qeval
Adversarial robustness Basic datasets MS MARCO document (Nguyen et al., 2016) 3.2M 370K 5,193 5,793
MS MARCO passage (Nguyen et al., 2016) 8.8M 500K 6,980 6,837
Clueweb09-B (Clarke et al., 2009) 50M 150
Natural Questions (Kwiatkowski et al., 2019) 21M 60K 8.8K 3.6K
TriviaQA (Joshi et al., 2017) 21M 60K 8.8K 11.3K
Expansion of datasets TREC DL19 (Craswell et al., 2020b) 43
TREC DL20 (Craswell et al., 2021) 54
TREC MB14 (Lin et al., 2013) 50
Off-the-shelf datasets ASRC (Raifer et al., 2017) 1,279 31
Q-MS MARCO (Liu et al., 2023d) 4,000
Q-Clueweb09 (Liu et al., 2023d) 292
DARA (Chen et al., 2023b) 164k 50k 3,490 3,489
Adaptation to a new corpus BEIR (Thakur et al., 2021) 18 corpus
OOD robustness Updates to a corpus CDI-MS (Chen et al., 2023d) 3.2M 370K 5,193 5,793
CDI-NQ (Chen et al., 2023d) 8.8M 500K 6,980 6,837
LL-LoTTE (Cai et al., 2023) 5.5M 16K 8.5K 8.6K
LL-MultiCPR (Cai et al., 2023) 3.0M 136K 15K 15K
Query variation DL-Typo (Zhuang and Zuccon, 2022) 60
noisy-MS MARCO (Campos et al., 2023) 5.6k
rewrite-MS MARCO (Campos et al., 2023) 5.6k
noisy-NQ (Campos et al., 2023) 2k
noisy-TQA (Campos et al., 2023) 3k
noisy-ORCAS (Campos et al., 2023) 20k
variations-ANTIQUE (Penha et al., 2022) 2k
variations-TREC19 (Penha et al., 2022) 430
Zhuang and Zuccon (2021) 41k
Unseen query type MS MARCO (Nguyen et al., 2016) 15k
L4 (Surdeanu et al., 2008) 10k
Appendix B The BestIR Benchmark
BestIR aims to provide a robustness evaluation benchmark for neural IR models. In order to address the comprehensive robustness challenge, we construct benchmarks mainly in terms of seven types of two aspects of robustness, i.e., adversarial robustness and OOD robustness.
For adversarial robustness, the datasets are usually used both for adversarial defense and adversarial attack tasks. There are three construction methodologies: (i) Basic datasets, which are the original IR datasets that are directly performed attacks and defenses on; (ii) Expansion of datasets, which are extensions of the original dataset to model unknown queries or new documents; and (iii) Off-the-shelf datasets, which are datasets customized for the task of adversarial attack or defense that can be used directly for evaluation.
For OOD robustness, datasets are used to evaluate the performance of the model under unseen documents and unseen queries, respectively. For each scenario, there are two types of evaluation perspectives. For unseen documents: (i) Adaptation to a new corpus consists mainly of IR datasets from different domains; and (ii) Updates to a corpus consist mainly of the same dataset sliced and diced based on factors such as time or randomly. For unseen queries: (i) Query variation consists of different variants of the same query intent, such as typos, changes in word order, and changes in the form of expression; and (ii) Unseen query type consists mainly of the different types of queries in a dataset.
Table 5 summarizes the statistics of the datasets provided in BestIR. BestIR is publicly available at https://github.com/Davion-Liu/BestIR. There are lots of datasets available within each aspect for robustness challenges and continue growing. We try to balance the assessment of each aspect of robustness to fully evaluate the model’s abilities. In the future, we will consider the issue of robustness in a broader sense and introduce datasets into BestIR. Meanwhile, researchers can also focus on observing specific aspects of robustness performance according to this categorization according to their concerns.